### v1

@@ Abstract
Despite the popularity of large language model (LLM) quantization for inference acceleration, significant uncertainty remains regarding the accuracy-performance trade-offs associated with various quantization formats. We present a comprehensive empirical study of quantized accuracy, evaluating popular quantization formats (FP8, INT8, INT4) across academic benchmarks and real-world tasks, on the entire Llama-3.1 model family. Additionally, our study examines the difference in text generated by quantized models versus their uncompressed counterparts. Beyond benchmarks, we also present a couple of quantization improvements which allowed us to obtain state-of-the-art accuracy recovery results. Our investigation, encompassing over 500,000 individual evaluations, yields several key findings: (1) FP8 weight and activation quantization (W8A8-FP) is lossless across all model scales, (2) INT8 weight and activation quantization (W8A8-INT), when properly tuned, incurs surprisingly low 1-3% accuracy degradation, and (3) INT4 weight-only quantization (W4A16-INT) is competitive with 8-bit integer weight and activation quantization. To address the question of the ”best” format for a given deployment environment, we conduct inference performance analysis using the popular open-source vLLM framework on various GPU architectures. We find that W4A16 offers the best cost-efficiency for synchronous deployments, and for asynchronous deployment on mid-tier GPUs. At the same time, W8A8 formats excel in asynchronous “continuous batching” deployment of mid- and large-size models on high-end GPUs. Our results provide a set of practical guidelines for deploying quantized LLMs across scales and performance requirements.
footnotetext: Correspondence to: $<$firstname$>$@neuralmagic.com.
@@ 1 Introduction
The massive computational costs of serving large language models have led to a significant amount of work on inference acceleration techniques, such as quantization (Frantar et al., 2022; Dettmers & Zettlemoyer, 2022; Lin et al., 2024a), speculative decoding (Chen et al., 2023; Leviathan et al., 2023), or model pruning (Xia et al., 2023; Muralidharan et al., 2024). Model quantization, i.e., reducing the bitwidth of weights, activations, or both, is arguably the standard approach to reducing memory and computational costs at deployment time. The critical trade-off in LLM quantization is between the speed/memory benefits and the accuracy of the resulting compressed model. Despite extensive research on quantization, few studies provide systematic benchmarks or practical guidelines on expected performance at different compression levels. This state of affairs is expressed in our fore-title, which uses a direct quote from the wave of user feedback regarding the quantized Llama-3.1-405B model (Dubey et al., 2024), initially suspected to drop significant accuracy (Zhang, 2024; Karpathy, 2024) and later shown to be near-lossless in LMSYS Arena user testing (Chiang et al., 2024). This uncertainty can be a significant barrier to adoption of quantized formats. In this paper, we systematically consider the following question:
@@
Due to the broad range of existing quantization techniques, we focus our study on formats with efficient computational support across different batching levels and low accuracy loss. This reduces the range of formats to be examined to 8-bit weights and activations (W8A8), in integer (INT) precision for NVIDIA Ampere and older devices, and floating-point (FP) precision for NVIDIA Hopper and Ada Lovelace, as well as 4-bit integer weights with 16-bit activations (W4A16-INT). These formats are efficiently supported in the vLLM inference engine (Kwon et al., 2023), which we use for performance experiments through the GuideLLM framework Neural Magic (2024).
To examine accuracy, we implemented a broad, automated evaluation suite, including leading academic benchmarks, namely the Open LLM Leaderboard V1 (Beeching et al., 2023), and its recent, more challenging, V2 version (Fourrier et al., 2024). We complement this examination with a set of real-world benchmarks, which focus on generative quality in an open-ended setup, specifically Arena-Hard-Auto-v0.1 Li et al. (2024b), HumanEval Chen et al. (2021), and HumanEval+ Liu et al. (2023a). We further extend these benchmarks with an in-depth text similarity analysis between outputs from uncompressed and quantized models. Finally, we examine the performance implications of deploying quantized LLMs. For this, we benchmark across all quantization formats in vLLM (version 0.6.1.post2) across three GPU types (A6000, A100, H100) in six latency-sensitive use cases and two throughput use cases. Our main findings are as follows:
@@
    W8A8-FP quantization is essentially lossless: Across all benchmarks, it preserves the uncompressed model’s average accuracy or falls within the evaluation’s margin of error. Remarkably, we achieve this via a simple and robust approach: dynamic per-token activation quantization and symmetric weight quantization via round-to-nearest assignment.
@@
    W8A8-INT quantization shows a tolerable accuracy loss, in the 1-3% per-task average range. This is achieved using dynamic activation quantization or the more complex SmoothQuant technique (Xiao et al., 2022); for weights, we use GPTQ (Frantar et al., 2022) with symmetric quantization. Our findings stand in contrast with prior work (Li et al., 2024a; Lee et al., 2024b), which reported large accuracy drops (more than 10% on average) in the same setting.
@@
    W4A16-INT quantization maintains consistently low accuracy loss, being competitive in terms of accuracy to W8A8-INT. Our findings are obtained for a simple variant of GPTQ, which we surprisingly find to be more accurate on real-world tasks than the more recent AWQ method Lin et al. (2024a).
@@
    On the text analysis side, we show that, in autoregressive text generation, larger quantized models tend to closely follow the original word choices and sentence structures of the baseline. At the same time, on smaller models, quantization introduces moderate structure variability relative to full precision, but still preserves semantic meaning.
@@
    For quantized inference performance, we show that W4A16-INT offers favorable performance in synchronous deployments. For asynchronous deployments, W4A16 is preferable for small models on mid-tier GPUs, and W8A8 formats are more effective for mid-to-large-size models on high-end GPUs. The most cost-effective quantization scheme depends on model size, hardware, use case (e.g., code completion vs. multi-turn chat), and deployment scenario (synchronous vs. asynchronous).
Overall, this work provides a first in-depth study of accuracy-vs-performance-vs-cost trade-offs for quantized models across various formats, algorithms, use cases, and hardware types. We hope it can serve as a guide for practitioners looking to deploy models efficiently, and as motivation to researchers interested in method improvements.
@@ 2.1 A Primer on Quantization
In the following, we provide some background on post-training quantization of LLM weights and activations. Early work on this topic has considered INT8 activation quantization and INT4/INT8 weight quantization (Dettmers et al., 2022; Yao et al., 2022; Park et al., 2022). A standard approach for weight quantization has been direct round-to-nearest (RTN) quantization over small groups: given a group of $g$ consecutive weights, viewed as a vector $\mathbf{x}\in\mathbb{R}^{g}$, $b$-bit RTN is defined as:
@@
where $\operatorname{rnd}$ rounds to the nearest integer, $z=z(\mathbf{x})=\min(\mathbf{x})$ is the “zero point” and $s=s(\mathbf{x})=(\max(\mathbf{x})-\min(\mathbf{x}))/(2^{b}-1)$ is the scale, taken as min-max. This early work raised two key issues: RTN on weights loses significant accuracy for INT4 precision, and RTN quantization of activations even to higher INT8 precision can be lossy due to the large “outlier features” in LLMs Dettmers et al. (2022).
Weight quantization. Towards the first issue, GPTQ (Frantar et al., 2022) improved upon RTN for scalar, per-weight, quantization by allowing weights to be adjusted via second-order updates computed over a sample of calibration data. Follow-up scalar quantization methods such as AWQ (Lin et al., 2024a), SqueezeLLM (Kim et al., 2023), OWQ (Lee et al., 2024a), and SpQR (Dettmers et al., 2023) implemented variants of outlier-aware quantization, where a small fraction of weights are effectively stored in higher precision. These methods can reach highly-accurate 4-bit quantization, but the latter three lack general kernel support.
More recently, several high-compression methods have been proposed, such as QuIP (Chee et al., 2023), QuIP# Tseng et al. (2024a), QTIP (Tseng et al., 2024b) AQLM (Egiazarian et al., 2024), and GPTVQ (van Baalen et al., 2024), targeting the 3-bit range or lower. These methods leverage much more complex representations, such as lattice/vector quantization, often paired with incoherence pre-processing of the weights. Despite high accuracy, these methods add overheads and lack efficient kernel support for batch sizes larger than 1, which is necessary for general deployments. Thus, our study does not currently consider these formats.
Activation quantization. Quantizing both activations and weights to 8-bit or 4-bit precisions allows direct utilization of existing hardware support for lower-bit operations. However, activation matrices are difficult to quantize effectively (Dettmers et al., 2022) due to the presence of outlier features — elements that can be up to 100 times larger than the average. Dettmers et al. (2022) extracted columns at runtime and quantized the rest to INT8, which is inefficient. SmoothQuant (Xiao et al., 2022) improves upon this by recognizing that outlier features are generally stable across the model and can be extracted utilizing a calibration set and folded during compilation.
Follow-ups to SmoothQuant investigate the possibility of W4A4 quantization (Ashkboos et al., 2023; 2024) or mixed-precision W4A8 (Lin et al., 2024b; Zhang et al., 2024), also considering quantizing the KV-cache. While promising, these approaches still lose significant accuracy and lack efficient support in high-performance inference frameworks. We, therefore, also leave their investigation for future work.
2.2 Efficient Inference With Quantized LLMs
Several performant LLM inference frameworks exist, such as HuggingFace TGI (HuggingFace, 2024), NVIDIA TensorRT-LLM (NVIDIA, 2023), and vLLM (Kwon et al., 2023). We focus our benchmarking efforts on vLLM, a popular open-source framework that addresses the specific challenges of serving large, autoregressive neural networks.
The backend architecture of vLLM combines efficient execution management with high-performance computation graphs that utilize pre-compilation, optimized operator fusion, and performant scheduling. Additionally, vLLM incorporates computational acceleration through customized kernels leveraging NVIDIA CUTLASS and hand-written GPU kernels. At the same time, vLLM enables broad hardware compatibility, supporting NVIDIA and AMD GPUs, TPUs, and CPUs. Several optimization techniques, including FlashAttention (Dao et al., 2022), FlashInfer (FlashInfer, 2023), PagedAttention (Kwon et al., 2023), and Marlin (Frantar et al., 2024) are also supported. This support and its open-source nature have led to vLLM’s adoption by industry and the open-source community. The framework integrates optimized CUTLASS kernels for W8A8 (FP and INT), and mixed-precision Marlin kernels for W4A16-INT.
2.3 Accuracy Evaluations of Quantized LLMs
There is a substantial amount of prior work on the evaluation of quantized LLMs, primarily focusing on accuracy trade-offs under various quantization schemes Yao et al. (2023); Liu et al. (2023b); Huang et al. (2024); Gong et al. (2024b); Li et al. (2024a); Gong et al. (2024a). However, much of this research has concentrated solely on academic benchmarks, which do not fully represent the real-world environments where LLMs are deployed. Furthermore, the fact that some of these works do not tune hyperparameters and algorithmic choices for the quantization algorithms can lead to incorrect conclusions regarding accuracy, as we demonstrate later in the experimental section. Specifically, we refer to the claim that 8-bit integer activation quantization causes significant accuracy drops Li et al. (2024a); Lee et al. (2024b), which we contradict experimentally.
The closest work to ours is by Lee et al. (2024b), which, like most prior studies, focuses solely on accuracy under quantization. However, we identify several points of divergence. First, although the authors claim to analyze models up to 405B in size, open-ended benchmarks are not included for this scale, and accuracy scores for the full-precision uncompressed model are missing even for academic benchmarks. Without baselines, it is difficult to draw meaningful conclusions on the impact of quantization at this scale. To address this, we enabled highly efficient multi-node evaluations for the 405B model, allowing comprehensive benchmarking in full precision across academic and real-world tasks.
The second divergence point concerns their claims that AWQ outperforms GPTQ in a 4-bit weight-only quantization setup. We attribute this finding to suboptimal hyperparameter choices in GPTQ quantization and provide a detailed comparison of the methods in Table 1 and Appendix A.2. Our results show that both methods perform similarly on academic benchmarks, while GPTQ demonstrates notable improvements over AWQ on real-world benchmarks, especially in coding tasks.
Third, we also challenge the prior conclusion that W8A8-INT performs substantially worse than W8A8-FP and W4A16-INT. When tuned carefully, W8A8-INT proves competitive, with only minor accuracy losses. For instance, while Lee et al. (2024b) reports a 10-point drop in accuracy for the W8A8-INT quantized 405B model on the Open LLM Leaderboard V2 benchmark relative to FP8 quantized one, our carefully optimized quantization parameters reduce this gap to just 0.7 points.
Unlike previous work, we also explore how quantization affects text generation quality. This analysis provides deeper insights into the impacts of various quantization schemes on sentence structure, word choice, and semantic fidelity in text generated by quantized LLMs.
Furthermore, we conclude our study by analyzing the performance impacts of deploying various quantization formats across a range of scenarios that represent real-world applications. This examination provides insights into how each quantization approach performs under practical deployment conditions and hardware configurations.
@@ 3.1 Datasets and Benchmarks
Our accuracy evaluation suite is designed to encompass a wide range of inference scenarios and use cases. It includes a diverse set of tasks to ensure a representative collection of topics relevant to practical deployments. Furthermore, the evaluation process is designed to be completely automated, as opposed to human-in-the-loop approaches, enabling streamlined experimentation. Overall, we group all of the benchmarks into three distinct categories.
Academic benchmarks such as Open LLM Leaderboard V1 and V2 Beeching et al. (2023); Fourrier et al. (2024), are essential for evaluating model improvements. They focus on structured tasks like question-answering and reasoning, providing consistent and reproducible scoring. However, they often lack alignment with real-world scenarios where semantics, variability, and context-awareness are important.
The Open LLM Leaderboard V1 consists of a diverse range of topics, including grade school math (GSM Cobbe et al. (2021)), world knowledge and reasoning (MMLU Hendrycks et al. (2020), ARC-Challenge Clark et al. (2018)), language understanding (Winogrande Sakaguchi et al. (2021), HellaSwag Zellers et al. (2019)), and truthfulness (TruthfulQA Lin et al. (2021)).
The Open LLM Leaderboard V2 builds on this with topics such as expert knowledge and complex reasoning (MMLU-Pro Wang et al. (2024), GPQA Rein et al. (2023), Big Bench Hard Suzgun et al. (2022)), multi-step reasoning (MuSR Sprague et al. (2024)), advanced math problems (MATH Level 5 Hendrycks et al. (2021)), and instruction following (IFEval Zhou et al. (2023)).
By evaluating with both Leaderboards, we cover a broad spectrum of topics. We employ log-likelihood (multiple-choice) and text-generation evaluations, allowing for stress testing of quantized LLMs under diverse conditions.
Real-world benchmarks, unlike academic benchmarks, test models in scenarios that mimic human usage, such as instruction following, chat, and code generation. These benchmarks include Arena-Hard-Auto-v0.1 Li et al. (2024b); Chiang et al. (2024); Li et al. (2024c), HumanEval Chen et al. (2021), and HumanEval+ Liu et al. (2023a), which offer a broader range of tasks with higher variation but better reflect real-world environments. The LMSYS Chatbot Arena Chiang et al. (2024) has established itself as a leading benchmark for LLMs, assessing how models align with human preferences. Arena-Hard-Auto-v0.1 is an automated extension of this benchmark, where an LLM judges responses to 500 complex prompts on various topics. It has demonstrated a strong correlation with human evaluations, achieving a state-of-the-art 89% agreement with human preference rankings Li et al. (2024c). This enables a reliable, fast, and automated evaluation of the model’s chat capabilities without requiring a human. In addition to chat-based interactions, LLMs are widely deployed as coding assistants. To evaluate the performance of quantized models in code generation, we tested them on HumanEval and its more challenging variant, HumanEval+. These benchmarks measure a model’s ability to generate correct and functional code based on programming problems, with HumanEval+ introducing more rigorous evaluation protocol.
Text similarity analysis evaluates how closely the generated outputs of quantized models align with those of their unquantized counterparts. While real-world benchmarks reflect practical usage scenarios for LLMs, their inherently open-ended format introduces notable variance in results. To address this, we extend our evaluation suite to include an analysis of output similarity under identical prompt requests. This allows us to get insights into how quantized models’ text generation differs from that of full-precision models. For this investigation, we employ metrics such as ROUGE Lin (2004), BERTScore Zhang et al. (2019), and Semantic Textual Similarity (STS) to assess both semantic and structural consistency, ensuring that the intended meaning and quality of generated text are preserved under quantization.
ROUGE-1 measures word-level (unigram) overlap between outputs from quantized and unquantized models, while ROUGE-L captures structural similarity by focusing on the longest common subsequence. BERTScore evaluates token-level contextual similarity, computed using the RoBERTa-large model’s contextual embeddings. STS assesses overall semantic similarity at the sentence level, using the Sentence Transformers framework Reimers & Gurevych (2019) built on the 6-layer MiniLM architecture Wang et al. (2020).
With this extensive evaluation framework, we ensured that deployment scenarios ranging from structured, research-driven tasks to open-ended, real-world applications were covered, providing a holistic view of the performance and capabilities of quantized LLMs.
3.2 Models, Formats, and Quantization Algorithms
We focus our evaluation on extensive testing of the recent Llama 3.1 series of models Dubey et al. (2024), which have gained significant traction in research and deployment contexts. To obtain a comprehensive understanding of trade-offs, we conduct all of our experiments on the instruction-tuned versions of all three available Llama 3.1 sizes (8B, 70B, and 405B), and for each of these, we focus on investigating three popular quantization formats with efficient inference support in vLLM: W8A8-FP, W8A8-INT, and W4A16-INT.

### v2

@@ Abstract
Quantization is a powerful tool for accelerating large language model (LLM) inference, but the accuracy-performance trade-offs across different formats remain unclear. In this paper, we conduct the most comprehensive empirical study to date, evaluating FP8, INT8, and INT4 quantization across academic benchmarks and real-world tasks on the entire Llama-3.1 model family. Through over 500,000 evaluations, our investigation yields several key findings: (1) FP8 (W8A8-FP) is effectively lossless across all model scales, (2) well-tuned INT8 (W8A8-INT) achieves surprisingly low (1-3%) accuracy degradation, and (3) INT4 weight-only (W4A16-INT) is more competitive than expected, rivaling 8-bit quantization. Further, we investigate the optimal quantization format for different deployments by analyzing inference performance through the popular vLLM framework. Our analysis provides clear deployment recommendations: W4A16 is the most cost-efficient for synchronous setups, while W8A8 dominates in asynchronous continuous batching. For mixed workloads, the optimal choice depends on the specific use case. Our findings offer practical, data-driven guidelines for deploying quantized LLMs at scale—ensuring the best balance between speed, efficiency, and accuracy.
“Give Me BF16 or Give Me Death”?
Accuracy-Performance Trade-Offs in LLM Quantization
Eldar Kurtic1,2, Alexandre Marques1, Shubhra Pandit1, Mark Kurtz1, Dan Alistarh1,2 1Red Hat AI, 2Institute of Science and Technology Austria Correspondence: ekurtic@redhat.com, dalistar@redhat.com
@@ 1 Introduction
The high computational cost of serving LLMs has driven extensive research into inference acceleration techniques, including quantization (Frantar et al., 2022; Dettmers and Zettlemoyer, 2022; Lin et al., 2024a), speculative decoding (Chen et al., 2023; Leviathan et al., 2023), and pruning (Xia et al., 2023; Muralidharan et al., 2024). Among these, quantization—reducing the bitwidth of weights, activations, or both—has emerged as the most widely used approach. However, its key challenge lies in balancing efficiency and accuracy.
Despite progress, systematic benchmarks and practical deployment guidelines remain scarce. This uncertainty has fueled speculation around quantized models, exemplified by the initial skepticism toward the Llama-3.1-405B quantized model release (Dubey et al., 2024), which was later found to be near-lossless in LMSYS Arena user evaluations (Chiang et al., 2024). To address this gap, we pose the following core question:
@@
In this study, we focus on widely supported, computationally efficient quantization formats. Specifically, we examine 8-bit weights and activations (W8A8), using integer (INT) precision for NVIDIA Ampere and older GPUs and floating-point (FP) precision for NVIDIA Hopper and Ada Lovelace. Additionally, we consider 4-bit integer weights with 16-bit activations (W4A16-INT), a competitive low-bit alternative. To evaluate accuracy, we implement a broad automated evaluation suite, spanning both academic and real-world benchmarks. Our academic benchmarks include Open LLM Leaderboard V1 (Beeching et al., 2023) and its more challenging V2 version (Fourrier et al., 2024), while real-world generative tasks are represented by Arena-Hard-Auto-v0.1 (Li et al., 2024b), HumanEval (Chen et al., 2021) and HumanEval+ (Liu et al., 2023a), and the long-context RULER benchmark (Hsieh et al., 2024). Beyond standard evaluations, we further analyze text similarity between outputs from uncompressed and quantized models to assess generative consistency.
Finally, we conduct an extensive inference performance study, benchmarking vLLM (Kwon et al., 2023) (version 0.6.4.post1) across three GPU architectures (A6000, A100, H100) in seven deployment scenarios. Our findings provide a comprehensive view of quantization’s trade-offs and offer practical recommendations for real-world LLM deployment. Our main findings are as follows:
@@
    W8A8-FP quantization is essentially lossless, preserving the uncompressed model’s accuracy across all benchmarks, often within the evaluation’s margin of error. This result is achieved with a simple yet robust approach: dynamic per-token activation quantization combined with symmetric weight quantization via round-to-nearest assignment.
@@
    W8A8-INT quantization exhibits only a modest accuracy degradation of 1–3% per task on average, far lower than the 10%+ drops reported in prior work (Li et al., 2024a; Lee et al., 2024b). This performance is enabled by dynamic activation quantization or SmoothQuant (Xiao et al., 2022), paired with GPTQ (Frantar et al., 2022) for symmetric weight quantization.
@@
    W4A16-INT quantization maintains consistently low accuracy loss, performing on par with W8A8-INT. Surprisingly, we show for the first time that a simple variant of GPTQ outperforms the more recent AWQ method (Lin et al., 2024a) on real-world tasks, challenging prior assumptions about low-bit quantization strategies.
@@
    Beyond accuracy, our text similarity analysis reveals that larger quantized models closely adhere to the word choices and sentence structures of their uncompressed counterparts in autoregressive text generation. In contrast, smaller quantized models introduce moderate variability in structure but still preserve semantic meaning.
@@
    In terms of performance, W4A16-INT is the most efficient choice for synchronous deployments, while W8A8 formats maximize throughput in asynchronous settings. The optimal quantization scheme depends on model size, hardware, and deployment needs—whether for latency-sensitive applications like code completion or high-throughput multi-turn chat.
Overall, this work provides the first in-depth study of accuracy vs. performance vs. cost trade-offs for quantized LLMs across formats, algorithms, use cases, and hardware types. We aim for these findings to serve as both a practical deployment guide and a strong and competitive foundation for future research on better quantization techniques.
@@ 2.1 A Primer on Quantization
Early work focused on INT8 activation quantization and INT4/INT8 weight quantization (Dettmers et al., 2022; Yao et al., 2022; Park et al., 2022). A common approach is round-to-nearest (RTN) over groups: given a group of $g$ consecutive weights as a vector $\mathbf{x}\in\mathbb{R}^{g}$, $b$-bit RTN is defined as:
@@
where $\operatorname{rnd}$ rounds to the nearest integer, $z(\mathbf{x})=\min(\mathbf{x})$ is the zero point, and $s(\mathbf{x})=(\max(\mathbf{x})-\min(\mathbf{x}))/(2^{b}-1)$ is the scale, computed using min-max normalization. However, RTN struggles at INT4 precision and suffers from lossy activation quantization even at INT8 (Dettmers et al., 2022).
Weight Quantization. To mitigate weight quantization errors, GPTQ (Frantar et al., 2022) introduced second-order weight adjustments using calibration data. Subsequent methods, including AWQ (Lin et al., 2024a), SqueezeLLM (Kim et al., 2023), OWQ (Lee et al., 2024a), and SpQR (Dettmers et al., 2023), incorporated outlier-aware quantization, storing a fraction of weights in higher precision to enable highly accurate 4-bit quantization. More recent high-compression techniques—QuIP (Chee et al., 2023), QuIP# Tseng et al. (2024a), QTIP (Tseng et al., 2024b), AQLM (Egiazarian et al., 2024), and GPTVQ (van Baalen et al., 2024)—target low-bitwidths using advanced representations such as vector quantization. Yet, these formats are inefficient for batch sizes larger than 1, limiting their practicality.
Activation Quantization. Quantizing both weights and activations enables low-bit hardware operations. Yet, activations are difficult to quantize due to outlier features—elements up to 100× larger than the average (Dettmers et al., 2022). Early attempts extracted outlier columns at runtime, but this is inefficient. SmoothQuant (Xiao et al., 2022) improves upon this by noticing that outliers are stable across the model and can be precomputed using a calibration set. Follow-up work explored W4A4 quantization (Ashkboos et al., 2023, 2024) and mixed-precision W4A8 (Lin et al., 2024b; Zhang et al., 2024), including KV-cache quantization. While promising, these methods still suffer accuracy loss and lack robust support in high-performance inference frameworks.
2.2 Related Work
A significant body of work has explored the accuracy trade-offs under different quantization schemes Yao et al. (2023); Liu et al. (2023b); Huang et al. (2024); Gong et al. (2024b); Li et al. (2024a); Gong et al. (2024a). However, much of this research relies primarily on academic benchmarks, which do not fully reflect real-world deployment scenarios. Additionally, the lack of hyperparameter tuning in some studies leads to misleading conclusions about accuracy, as we demonstrate in our experiments. We challenge the claim that 8-bit integer activation quantization causes substantial accuracy degradation Li et al. (2024a); Lee et al. (2024b), providing vast evidence to the contrary.
The closest work to ours is by Lee et al. (2024b), which, like most prior studies, focuses on quantization accuracy, but overlooks key factors. First, while the authors claim to analyze models up to 405B parameters, they omit open-ended benchmarks at this scale and fail to report full-precision baselines even for academic tasks. Without these references, the impact of quantization remains unclear. To address this, we enable efficient multi-node evaluations for the 405B model, conducting a comprehensive accuracy analysis in both academic and real-world settings. Second, Lee et al. (2024b) asserts that AWQ outperforms GPTQ in a 4-bit weight-only quantization setup. We correct this claim, and attribute it to suboptimal hyperparameter choices. Our comparative analysis (Table 1 and Appendix A.2) shows that while both methods perform similarly on academic benchmarks, GPTQ exhibits notable gains over AWQ in real-world tasks, particularly coding.
Third, we refute the conclusion that W8A8-INT is significantly inferior to W8A8-FP and W4A16-INT. With proper tuning, W8A8-INT achieves competitive accuracy, with only minor losses. For example, while Lee et al. (2024b) reports a 10-point accuracy drop for W8A8-INT quantized 405B models on the Open LLM Leaderboard V2 compared to FP8, our approach reduces this to just 0.7 points.
@@ 3.1 Datasets and Benchmarks
We categorize benchmarks into three groups: academic, real-world, and text similarity analysis.
1. Academic benchmarks, such as Open LLM Leaderboard V1 and V2 Beeching et al. (2023); Fourrier et al. (2024), provide structured evaluations for question-answering and reasoning tasks. While widely used for benchmarking, they lack alignment with real-world scenarios involving semantics, variability, and context-awareness. Leaderboard V1 includes tasks like GSM for grade school math Cobbe et al. (2021), MMLU and ARC-Challenge for world knowledge and reasoning Hendrycks et al. (2020); Clark et al. (2018), Winogrande and HellaSwag for language understanding Sakaguchi et al. (2021); Zellers et al. (2019), and TruthfulQA for factual correctness Lin et al. (2021). Leaderboard V2 extends this with expert knowledge benchmarks such as MMLU-Pro Wang et al. (2024), GPQA Rein et al. (2023), and Big Bench Hard Suzgun et al. (2022), as well as multi-step reasoning (MuSR Sprague et al. (2024)), advanced math (MATH Level 5 Hendrycks et al. (2021)), and instruction following (IFEval Zhou et al. (2023)). By evaluating across both leaderboards, we capture a broad spectrum of reasoning and knowledge domains, using both log-likelihood and text-generation evaluations to stress-test quantized models.
@@
2. Real-world benchmarks evaluate models in practical scenarios such as instruction following, chat, long-context, and code generation. Arena-Hard-Auto-v0.1 (Li et al., 2024b; Chiang et al., 2024; Li et al., 2024c) automates LMSYS Chatbot Arena (Chiang et al., 2024) evaluations, using an LLM to judge responses to 500 complex prompts, achieving an 89% agreement with human rankings (Li et al., 2024c). This allows rapid and scalable assessment of chat capabilities without human intervention. For code generation, we evaluate models on HumanEval (Chen et al., 2021) and its extension HumanEval+ (Liu et al., 2023a), which test the ability to generate correct and functional code. Finally, we conduct long-context evaluations via the rigorous RULER benchmark (Hsieh et al., 2024) which consists of retrieval, multi-hop tracing, information aggregation, and question answering evaluations at sequence lengths from 4k to 128k.
3. Our text similarity analysis benchmark assesses how closely quantized models’ outputs align with their full-precision counterparts. While real-world benchmarks reflect practical usage, their open-ended nature introduces variability, making direct accuracy comparisons challenging. To mitigate this, we analyze output similarity under identical prompts using ROUGE (Lin, 2004), BERTScore (Zhang et al., 2019), and Semantic Textual Similarity (STS) (Reimers and Gurevych, 2019). ROUGE-1 measures unigram overlap, while ROUGE-L captures structural similarity through the longest common subsequence. BERTScore computes token-level contextual similarity using RoBERTa-large embeddings, and STS assesses semantic alignment at the sentence level via Sentence Transformers built on MiniLM (Wang et al., 2020).
3.2 Models, Formats, and Algorithms
We evaluate using the highly-popular Llama 3.1 model series Dubey et al. (2024). To assess quantization trade-offs, we conduct experiments on the instruction-tuned versions of all available sizes (8B, 70B, and 405B). For each, we examine the three main formats with kernel support in vLLM: W8A8-FP, W8A8-INT, and W4A16-INT.